UT 20120404 AO#2 Commissioning

Observer: ARiccardi, LBusoni, FQuiros-Pacheco, CArcidiacono (LBTO)
Telescope Operator: GBechetti (LBTO)
Additional AO Support: DMiller
Telescope Support: MMidkiff(LBTO)
Instrument Support: DOfficer (LBTO)

Summary

A BCU board in Adsec SX failed to communicate after about an hour on sky. This was the same board that was replaced a couple weeks ago, but the problem was different. Armando said tonights problem was the same as the failure of a BCU in crate 5 of Adsec DX a few months ago.

After getting back on sky, the Arcetri team completed the following:
  • Close loop on a bright star to confirm the performance is the same as a month ago. Performance confirmed, 90% strehl in H.
  • Close loop on a bright star, but with Fernado's synthetic reconstructor, to measure performance.
  • Determine best focus for IRTC J and H filters

Details

02:54 Acquire preset to set pointing

03:00 Active preset to BS star. Collimate. Add 3000+ nm

03:10 ACE preset to AO366

03:15 RunAO failed because of incorrect parameter. In the presetAO sent above, the AOS had set the "WFS spec" (see AOS command GUI) to None. It should be FLAOWFS. I set this on the AOS command gui and resend the presetAO (from the GUI). Luca needs to fix this. He did: IssueTrak# 3947. Will be in next build.

03:23 RunAO successful. Closed AO Loop on a bright star AO366.

03:31 Shell ripped. Masterdiagnostic process die. Armando have to power off Adsec and power on. Geno re-home hexapod

03:56 Shell ripped again. Armando power cycle. Find that mbcu1sx is not responding to a ping. This is the same board that had to be replaced a few weeks ago. IT#3997.

04:22 Armando power off the SX unit and let it sit for 5 minutes. Power on and now mbcu1sx is communicating. Set Shell and re init the hexapod.

06:25 Armando and crew have tested fiber connections and put in new fiber and the problem remains. The board will have to be switched out. Armando will have to modify a board (ie change cold plate on a board made for right side of crate to fit in left side of crate). This will not take too long, but then will have to get it installed and burn the new IP address (will need sw from Marco to install IP). They found that carrier for the ethernet communication from the BCU to the ethernet switch was working properly (this was the problem a couple weeks ago). Thus, there is some problem getting communication from the ethernet board to the BCU board.

08:15 The bad board has been removed and Armando is now moving the cold plate from the old to the new BCU board.

11:12 Secondary is operational and loop is closed on-sky

11:30 DIMM Seeing 0.8. 90% strehl in H

12:40 Close

-- DougMiller - 04 Apr 2012
Topic revision: r4 - 04 Apr 2012, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback